# Unifews: Unified Entry-Wise Sparsification for Efficient Graph Neural Network

### v1

@@ 5.2. Decoupled Case: Graph Sparsification
We briefly explain the statement as below. Recall that common GNN learning can be expressed by the graph smoothing process Definition 3.1. Such optimization problem can be iteratively solved by employing a gradient descent scheme (Ma et al., 2021), where each iteration derives the $l$-th hop of graph propagation as in Eq. 4:
(17) $\bm{p}_{(l+1)}=\bm{p}_{(l)}-\frac{b}{2}\cdot\left.\frac{\partial\mathcal{L}}{\partial\bm{p}}\right|_{\bm{p}=\bm{p}_{(l)}}=(1-b)\bm{p}_{(l)}-bc\bm{L}\bm{p}_{(l)}+b\bm{x}.$
@@
where $b/2$ is the step size and initially there is $\bm{p}_{(0)}=\bm{x}$. Eq. 17 is expressive to represent various propagation operations in decoupled GNN models. For example, APPNP (Klicpera et al., 2019) can be achieved by letting $b=\alpha$ and $c=(1-\alpha)/\alpha$, while SGC (Wu et al., 2019) is the edge case with only the graph regularization term.
@@
Now consider the layer-wise graph sparsification under Eq. 17 updates. For the initial state, there is $\bm{p}_{(0)}=\hat{\bm{p}}_{(0)}=\bm{x}$. If in the $l$-th hop, Eq. 8 edge sparsification is applied to the graph $\bm{L}$, then the approximation gap is:
(18) $\bm{p}_{(l+1)}-\hat{\bm{p}}_{(l+1)}=(1-b)(\bm{p}_{(l)}-\hat{\bm{p}}_{(l)})+bc(\hat{\bm{L}}\hat{\bm{p}}_{(l)}-\bm{L}\bm{p}_{(l)}).$
@@
To demonstrate that $\hat{\bm{p}}_{(l+1)}$ is an $\epsilon$-approximation, we use induction by assuming $\hat{\bm{p}}_{(l)}-\bm{p}_{(l)}=\bm{\Delta}_{p}$, $\|\bm{\Delta}_{p}/\bm{p}_{(l)}\|\sim O(\epsilon)$. Then:
@@
where the inequalities follows from the property of matrix spectral norm and Eq. 16. Hence, the relative error of approximate representation $\hat{\bm{p}}_{(l+1)}$ is constrained by $O(\epsilon)$.
An empirical evaluation of Proposition 5.4 is shown in Figure 2(b), which validates that the approximation error is decided by the sparsification threshold as well as propagation hops. Additionally, we would like to remark that the “error” induced by Unifews is not necessarily harmful. Examining Eq. 4, GNNs aim to reach the optimization goal through multi-layer propagation, in which the first loss term representing the distance to original input attributes is increased, and graph smoothing is emphasized by a decreasing second term. Excessive propagation can lead to a decline in performance known as the over-smoothing issue (Li et al., 2018; Zhu et al., 2024), where the graph regularization is dominant and meaningful node identity is lost. Conversely, by eliminating a portion of graph diffusion, Unifews allows for alleviating the issue for deeper layers. We showcase its effect by depicting the embedding difference to optimization convergence in Figure 2(c). With an increased number of hops, the output embedding tends towards an over-smoothed state. However, a stronger sparsification prevents the rapid smoothing and hereby contributes to better graph learning performance.
5.3. Iterative Case: Joint Sparsification
To apply the approximation analysis to iterative GNNs, we first extend the analysis to multi-feature input matrix. Since graph operations among feature dimensions are mutually independent, results in Section 5.2 are still valid in their matrix forms. In iterative models, the gradient update similar to Eq. 17 is instead employed to the representation matrix $\bm{H}_{(l)}$ and derives each layer as:
@@ 5.3. Iterative Case: Joint Sparsification
As detailed in (Ma et al., 2021; Zhu et al., 2021), the update scheme is able to describe an array of iterative GNNs. For instance, when $\bm{H}_{(l)}=\bm{X}$ and $b=1/c$, it yields the GCN propagation $\bm{P}_{(l)}=\tilde{\bm{A}}\bm{H}_{(l)}$.
To interpret the effect of Unifews sparsification, we first investigate the entry-wise graph pruning in Algorithm 2 and its outcome, i.e., the embedding matrix $\hat{\bm{P}}_{(l)}$. For simplicity, we assume the sparsification is only employed upon the $(l+1)$-th layer and $\hat{\bm{H}}_{(l)}=\bm{H}_{(l)}$. Invoking Eq. 16 and the fact that $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{2}\|\bm{B}\|_{F}$, the margin of approximate embeddings can be written as:
@@
Then, consider the weight pruning Eq. 12. The entry-wise error is a composition of joint embedding and weight approximation:
@@
Let $\bm{M}_{(l)}=\bm{W}_{(l)}-\hat{\bm{W}}_{(l)}$. Recalling Eq. 1, the total difference on linear transformation $\bm{H}^{\prime}=\bm{P}_{(l)}\bm{W}_{(l)}$ is built up by:
$\small\hat{\bm{H}}^{\prime}-\bm{H}^{\prime}=\sum_{i,j=1}^{f}\big{(}\hat{\bm{P}}_{(l)}[:,j]-\bm{P}_{(l)}[:,j]\big{)}\bm{W}_{(l)}[j,i]+\hat{\bm{P}}_{(l)}[:,j]\bm{M}_{(l)}[j,i].\vspace{-3pt}$
@@
and the second term adheres to weight sparsification as per Eq. 12:
@@
$\|\hat{\bm{H}}_{(l+1)}-\bm{H}_{(l+1)}\|_{F}\leq\ell_{\sigma}bc\epsilon\|\bm{H}_{(l)}\|_{F}\|\bm{W}\|_{F}+\ell_{\sigma}q_{w}\delta_{w},$
where $\ell_{\sigma}$ is the Lipschitz constant representing the non-linearity of activation function $\sigma$ (Virmaux and Scaman, 2018).
@@
Table 2. Statistics of graph datasets. $f$ and $N_{c}$ are the numbers of input attributes and label classes, respectively. “Train” column is the percentage of nodes for training.
Dataset Nodes $n$ Edges $m$ $f$ $N_{c}$ Train cora $2,485$ $12,623$ $1433$ $7$ $50\%$ citeseer $3,327$ $9,228$ $3703$ $6$ $50\%$ pubmed $19,717$ $88,648$ $500$ $3$ $50\%$ arxiv $169,343$ $2,315,598$ $128$ $40$ $54\%$ products $2,400,608$ $123,718,024$ $100$ $47$ $8\%$ papers100m $111,059,956$ $3,228,124,712$ $128$ $172$ $78\%$
@@
Figure 4. Accuracy results of Unifews and baseline methods applied on iterative backbone models GCN and GAT over three small datasets. Columns of “Edge” and “Weight Sparsity” present the average results of models with solely edge and weight sparsification, respectively. Black dashed lines are the performance of backbone models with full graph and weights.
6. Experimental Evaluation
We implement Unifews for sparsifying various GNN architectures and evaluate its performance against strong competitors. In Sections 6.2 and 6.3, we separately conduct comparison for iterative and decoupled models, respectively highlighting the accuracy and efficiency improvement of our approach. The joint pruning and related parameters are discussed in Sections 6.4 and 6.5.
6.1. Experiment Settings
[…]
Dataset. We adopt 3 representative small-scale datasets: cora, citeseer, and pubmed (Kipf and Welling, 2017) as well as 3 large-scale ones: arxiv, products, papers100m (Hu et al., 2020). Statistics of the datasets can be found in Table 2, where we follow the common settings including the graph preparing and training set splitting pipeline in these benchmarks. In the table, we incorporate self-loop edges and count undirected edges twice to better reflect the propagation overhead.
[…]
Metrics. We uniformly utilize transductive node classification accuracy as the evaluation metric of model prediction. To precisely observe and compare the number of operations in GNN learning, the efficiency is assessed by computation time and floating-point operations (FLOPs), and $1$\mathrm{F}\mathrm{L}\mathrm{O}\mathrm{P}\mathrm{s}$\approx 2$\mathrm{M}\mathrm{A}\mathrm{C}\mathrm{s}$$. The graph and weight sparsities are calculated as portions of pruned entries compared to the original matrices. For layer-dependent methods, the average sparsity across all layers is used. Evaluation are conducted on a server with 32 Intel Xeon CPUs (2.4GHz), an Nvidia A30 (24GB memory) GPU, and 512GB RAM.
[…]
Backbone Models. Since Unifews is adaptable to a range of GNN architectures, we select 4 classic GNNs as our backbones, i.e., subjects of Unifews and baseline sparsification methods. We divide them into two categories based on GNN designs introduced in Section 3.1. Iterative backbone models include:
  • GCN (Kipf and Welling, 2017) is the representative message-passing GNN with a diffusion matrix $\bm{T}=\tilde{\bm{A}}$ across all layers.
Table 3. Evaluation results of $\eta_{a}=50\%$ graph sparsification on decoupled models. Unifews is applicable to both SGC and APPNP backbones, while two baseline methods are only available on the former one. “Acc” and “Time” are the prediction accuracy (%) and propagation time (in seconds), respectively. “FLOPs” separately presents the operational complexity (in $\mathrm{G}\mathrm{M}\mathrm{A}\mathrm{C}\mathrm{s}$) for propagation and transformation on all nodes. “$\mathrm{O}\mathrm{O}\mathrm{M}$” stands for out of memory error. “Improvement” is the comparison between Unifews and the corresponding backbone model, where improved accuracy is marked in bold fonts.
Dataset cora citeseer pubmed arxiv products papers100m Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs SGC 85.8 0.13 0.36+2.2 67.7 0.44 0.68+2.8 83.0 0.36 0.89+2.3 68.8 3.9 5.9+15.0 79.1 289.6 247.4+434.4 63.3 19212 17.7+253.6 +NDLS 80.3 1362 0.19+2.7 64.7 1940 0.33+4.3 77.0 4717 0.42+2.0 (OOM) (OOM) (OOM) +NIGCN 84.2 0.45 0.22+3.0 70.4 0.47 0.28+2.1 85.0 9.2 0.44+2.0 63.7 87.6 15.6+14.7 77.9 1026 137.2+182.0 53.7 1770 110.7+238.6 +Unifews (Ours) 86.0 0.10 0.18+1.2 73.0 0.26 0.35+1.7 83.0 0.24 0.47+2.0 69.4 1.5 3.0+13.3 78.5 203.1 124.0+186.9 63.1 192.4 5.3+143.8 Improvement 0.2 1.3× 2.0×, 1.9× 5.3 1.7× 2.0×, 1.7× 0.0 1.5× 1.9×, 1.2× 0.6 2.6× 2.0×, 1.1× -0.5 1.4× 2.0×, 2.3× -0.2 99.8× 3.3×, 1.8× APPNP 86.2 0.15 0.36+2.2 71.6 0.43 0.68+2.8 87.6 0.33 0.89+2.3 64.8 2.6 5.9+20.9 72.5 248.5 247.4+269.4 60.9 15305 17.7+247.7 +Unifews (Ours) 86.5 0.08 0.18+1.8 73.7 0.21 0.31+2.3 88.0 0.26 0.43+1.8 65.0 0.93 3.0+15.0 76.9 58.5 31.7+186.9 62.8 178.5 8.9+241.8 Improvement 0.4 1.8× 2.0×, 1.2× 2.1 2.0× 2.2×, 1.2× 0.4 1.3× 2.1×, 1.3× 0.2 2.8× 1.9×, 1.4× 4.5 4.2× 7.8×, 1.4× 1.9 85.7× 2.0×, 1.0×
[…]
Baseline Methods. Our selection of baselines is dependent on their applicable backbones. We mostly utilize their public source codes and retain original implementations. For iterative models, we consider state of the arts in graph and joint compression, which are methods with the capability to produced sparsified graphs with smaller edge sets for both GNN training and inference:
  • Random refers to the sparsification method that removes entries with uniform probability based on the specified sparsity.
[…]
Hyperparameters. We commonly utilize graph normalization $r=0.5$, model layer depth $L=2$, and layer width $f_{hidden}=512$. For decoupled models, the number of propagation hops is $20$. We employ full-batch and mini-batch training for iterative and decoupled methods, respectively. The total number of training epochs is uniformly $200$ for all models, including pre-training or fine-tuning process in applicable methods. The batch size is $512$ for small datasets and $16384$ for large ones. We tune the edge and weight sparsity of evaluated models across their entire available ranges, and the pruning ratio of Unifews is controlled by adjusting $\delta_{a}$ and $\delta_{w}$.
(a) Edge Pruning on cora
(b) Edge Pruning on citeseer
Figure 5. Accuracy results of Unifews and baseline methods applying graph sparsification on decoupled models over cora and citeseer. Unifews is employed with solely edge removal and $\eta_{w}=0\%$. Black and brown dashed lines are the performance of backbone SGC and APPNP models with full graph and weights, respectively.
6.2. Accuracy on Iterative GNN
[…]
Overview. In this section, we first focus on separately applying one part of the sparsification, i.e., either edge or weight pruning, and compare the model performance. Figure 4 presents accuracy results of iterative backbones and compression methods over cora, citeseer, and pubmed datasets. GEBT encounters out of memory error on pubmed. For larger graphs, most of the baselines suffer from the out of memory error due to the expense of trainable adjacency matrix design and full-graph training process, hence the results are not presented. In contrast, we demonstrate the ability of Unifews for boosting models on large graphs in later sections.
As an overview, for both graph and weight sparsification, our Unifews approach successfully removes up to 90%-95% of entries while consistently attaining comparable or better prediction accuracy. Considering that the number of operations is proportional to density, it indicates that Unifews is capable of achieving 10-20$\times$ computational reduction for GNN training and inference.
(a) Edge Pruning with $\eta_{w}=30\%$
(b) Weight Pruning with $\eta_{a}=30\%$
Figure 6. Accuracy results of joint Unifews and baseline methods on GCN over cora. Columns of “Edge” and “Weight Sparsity” respectively present the results of models with varying edge and weight sparsification, while fixing the other sparsity at $30\%$. Black dashed lines are the performance of backbone GCN with full graph and weights.
(a) GCN on arxiv
(b) GAT on arxiv
(c) SGC on arxiv
(d) SGC on products
Figure 7. Composition of full-graph inference FLOPs with respect to propagation and transformation operations, under varying Unifews edge and weight thresholds. Iterative models: (a) GCN and (b) GAT over arxiv. Decoupled models: SGC over (c) arxiv and (d) products datasets. For better presentation, model width is set to 64 in this experiment.
[…]
Edge Sparsification. As specifically shown in Figures 3(a) and 3(c), Unifews outperforms state-of-the-art graph and joint compression approaches in most backbone-dataset combinations. Typically, for relatively small ratios $\eta_{a}<80\%$, models with Unifews pruning achieve comparable or exceeding accuracy, aligning with our approximation analysis. For higher sparsity, Unifews benefits from the skip connection design, which carries essential information of node identity and therefore retains accuracy no worse than the trivial transformation. On the contrary, most of the competitors experience significant accuracy drop on one or more datasets. On cora and citeseer where graph topology is important, an aggressive pruning causes noticeable performance loss, which limits their application to highly-sparsified graphs. CGP and DSpar exhibit poor utility on the GAT backbone since its variable connections are more vulnerable to removal. Additionally, the comparison with random sparsification indicates that, the entry-wise scheme is particularly effective for small ratios, as the randomized pruning even surpasses dedicated methods for some circumstances. For high edge sparsity, the Unifews preservation of dominant edges and identity mapping is critical to accuracy, which validates the advantage of our design.
[…]
Weight Sparsification. Figures 3(b) and 3(d) display the efficacy of network pruning. For all combinations of models and graphs, Unifews achieves top-tier accuracy with a wide range of weight sparsity. For small ratios, the model produces up to $3\%$ accuracy gain over the bare GNN, resembling the Occam’s hill phenomenon for general neural network compression (Hoefler et al., 2021). In short, it is hypothesized that pruning can be regarded as a weight regularization process that improves model generality and alleviates over-fitting, and thus favorable for network transformation. Most evaluated baselines also maintain low error rate compared to their performance in graph pruning, indicating that the GNN weights are relatively redundant and are suitable for substantial compression.
6.3. Efficiency on Decoupled GNN
[…]
FLOPs. Decoupled models are born with the capability of large-scale task. Hence, we evaluate the performance by Table 3, particularly focusing on the efficiency aspect of graph propagation. We utilize a representative graph sparsification ratio $\eta_{a}=50\%$, and compare the running time as well as FLOPs for decoupled models. Evaluation on the propagation FLOPs implies that, Unifews is effective in producing operational reduction proportional to the sparsity, i.e., saving $50\%$ computation FLOPs. Contrarily, NDLS suffers from out of memory error on large datasets due to its complex calculation and implementation, while NIGCN does not guarantee decreased computation because of its expansionary propagation design. Table 3 also presents the transformation FLOPs for reference, where the sparsified embedding of Unifews also benefits the downstream computation. We nonetheless remark that, although the FLOPs value for transformation stage appears to be on the same level with propagation, in practice it can be efficiently processed utilizing parallel computation on devices such as GPU (Liu et al., 2020; Peng et al., 2022; Fang et al., 2023a). Hence the GNN scalability bottleneck, especially on large datasets, is still the graph propagation. In this context, the ability of Unifews in reducing propagation operations is of unique importance.
[…]
Runtime Analysis. Within each method in Table 3, the propagation time is correlated with its FLOPs, and escalates rapidly with the data size due to the nature of graph operations. Comparison across methods demonstrates that Unifews excels in efficiently conducting graph propagation with regard to both FLOPs and execution time. It realizes significant acceleration on large graphs by favorably reducing the graph scale across all layers, which benefits the system workload such as better entry-level access. The papers100m result highlights the superiority of Unifews with 85100×85-100$\times$85 - 100 × improvement over the backbone and 9×9$\times$9 × speed-up over NIGCN. We examine the effectiveness of decoupled propagation in Figure 5 as well. Similarly, Unifews is able to preserve remarkable accuracy even with high sparsity. On citeseer, it also raises SGC accuracy by $2\%$ through mitigating the over-smoothing issue.
(a) Message $\bm{\tau}_{(0)}$
(b) Representation $\bm{H}_{(1)}$
Figure 8. Density of intermediate results during joint edge and weight sparsification at varying threshold values, evaluated by the ratio of matrix norm to the unpruned one on a 2-layer GCN over cora. (a) Density of edge-wise message $\|\hat{\bm{\tau}}_{(0)}\|/\|\bm{\tau}_{(0)}\|$ relative to the raw one in the first layer $l=0$. (b) Density of node-wise representation $\|\hat{\bm{H}}_{(1)}\|_{F}/\|\bm{H}_{(1)}\|_{F}$ relative to the raw one output by the first layer.

### v2

@@ Appendix B Detailed Proof and Theoretical Analysis
Recall that common GNN learning can be expressed by the graph smoothing process Definition 3.1. Such optimization problem can be iteratively solved by employing a gradient descent scheme (Ma et al., 2021), where each iteration derives the $l$-th hop of graph propagation as in Equation 1:
$\bm{p}_{(l+1)}=\bm{p}_{(l)}-\frac{b}{2}\cdot\left.\frac{\partial\mathcal{L}}{\partial\bm{p}}\right|_{\bm{p}=\bm{p}_{(l)}}=(1-b)\bm{p}_{(l)}-bc\bm{L}\bm{p}_{(l)}+b\bm{x}.$ (14)
@@
where $b/2$ is the step size and initially there is $\bm{p}_{(0)}=\bm{x}$. Equation 14 is expressive to represent various propagation operations in decoupled GNN models. For example, APPNP (Klicpera et al., 2019) can be achieved by letting $b=\alpha$ and $c=(1-\alpha)/\alpha$, while SGC (Wu et al., 2019) is the edge case with only the graph regularization term.
@@
Now consider the layer-wise graph sparsification under Equation 14 updates. For the initial state, there is $\bm{p}_{(0)}=\hat{\bm{p}}_{(0)}=\bm{x}$. If in the $l$-th hop, Equation 5 edge sparsification is applied to the graph $\bm{L}$, then the approximation gap is:
$\bm{p}_{(l+1)}-\hat{\bm{p}}_{(l+1)}=(1-b)(\bm{p}_{(l)}-\hat{\bm{p}}_{(l)})+bc(\hat{\bm{L}}\hat{\bm{p}}_{(l)}-\bm{L}\bm{p}_{(l)}).$ (15)
@@
To demonstrate that $\hat{\bm{p}}_{(l+1)}$ is an $\epsilon$-approximation, we use induction by assuming $\hat{\bm{p}}_{(l)}-\bm{p}_{(l)}=\bm{\Delta}_{p}$, $\|\bm{\Delta}_{p}/\bm{p}_{(l)}\|\sim O(\epsilon)$. Then the bound for approximation follows:
@@
where the inequalities follows from the property of matrix spectral norm and Equation 10. Hence, the relative error of approximate representation $\hat{\bm{p}}_{(l+1)}$ is constrained by $O(\epsilon)$.
(a) Distribution of Entries
(b) Approximate Embedding Error
(c) Distance to Over-smoothed Embedding
Figure 9: Empirical evaluation on the SGC graph propagation and the effect of Unifews on cora dataset. (a) Distribution of entries in the diffusion matrix $\bm{T}$ and the message $\bm{\tau}_{(0)}$ at hop $l=0$. (b) The relative error margin $\|\hat{\bm{P}}_{(L)}-\bm{P}_{(L)}\|_{F}/\|\bm{P}_{(L)}\|_{F}$ of approximation to the raw embedding against the strength of Unifews and propagation hops. (c) The relative distance $\|\hat{\bm{P}}_{(L)}-\bm{P}^{\ast}\|_{F}/\|\bm{P}^{\ast}\|_{F}$ to the converged embedding against the strength of Unifews and propagation hops.
B.5 Proof Sketch of Proposition 4.4
To apply the approximation analysis to iterative GNNs, we first extend the analysis to multi-feature input matrix. When there are $f$ input vectors, i.e., the input matrix is $\bm{X}\in\mathbb{R}^{n\times f}$, then the matrix form of graph Laplacian smoothing corresponding to Equation 1 is:
$\bm{P}^{\ast}=\operatorname*{arg\,min}_{\bm{P}}\|\bm{P}-\bm{X}\|_{F}^{2}+c\cdot\operatorname{tr}(\bm{P}^{\top}\bm{L}\bm{P}),$ (16)
where $\|\cdot\|_{F}$ is the matrix Frobenius norm and the closed-form solution is $\bm{P}^{\ast}=(\bm{I}+c\bm{L})^{-1}\bm{X}$.
Since graph operations among feature dimensions are mutually independent, conclusion from Theorem 4.2 and Proposition 4.3 are still valid in their matrix forms. In iterative models, the gradient update similar to Equation 14 is instead employed to the representation matrix $\bm{H}_{(l)}$ and derives each layer as:
@@
As detailed in (Ma et al., 2021; Zhu et al., 2021), the update scheme above is able to describe an array of iterative GNNs. For instance, when $\bm{H}_{(l)}=\bm{X}$ and $b=1/c$, it yields the GCN propagation $\bm{P}_{(l)}=\tilde{\bm{A}}\bm{H}_{(l)}$. GAT convolution can be imitated by a node-wise $c$ depending on the attention values (Veličković et al., 2018; Brody et al., 2022).
To interpret the effect of Unifews sparsification, we first investigate the entry-wise graph pruning in Algorithm 2 and its outcome, i.e., the embedding matrix $\hat{\bm{P}}_{(l)}$. For simplicity, we assume the sparsification is only employed upon the $(l+1)$-th layer and $\hat{\bm{H}}_{(l)}=\bm{H}_{(l)}$. Invoking Equation 10 and the fact that $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{2}\|\bm{B}\|_{F}$, the margin of approximate embeddings can be written as:
@@
Then, consider the weight pruning Equation 9. The entry-wise error is a composition of joint embedding and weight approximation:
@@
Let $\bm{M}_{(l)}=\bm{W}_{(l)}-\hat{\bm{W}}_{(l)}$. Recalling the iterative update scheme Equation 8, the total difference on linear transformation $\bm{H}^{\prime}=\bm{P}_{(l)}\bm{W}_{(l)}$ is built up by:
$\small\hat{\bm{H}}^{\prime}-\bm{H}^{\prime}=\sum_{i,j=1}^{f}\big{(}\hat{\bm{P}}_{(l)}[:,j]-\bm{P}_{(l)}[:,j]\big{)}\bm{W}_{(l)}[j,i]+\hat{\bm{P}}_{(l)}[:,j]\bm{M}_{(l)}[j,i].$
@@
and the second term adheres to weight sparsification as per Equation 9:
@@
$\|\hat{\bm{H}}_{(l+1)}-\bm{H}_{(l+1)}\|_{F}\leq\ell_{\sigma}bc\epsilon\|\bm{H}_{(l)}\|_{F}\|\bm{W}\|_{F}+\ell_{\sigma}q_{w}\delta_{w},$ (17)
where $\ell_{\sigma}$ is the Lipschitz constant representing the non-linearity of activation function $\sigma$ (Virmaux & Scaman, 2018). The above equation corresponds to the one in Proposition 4.4.
@@
Table 3: Statistics of graph datasets. $f$ and $N_{c}$ are the numbers of input attributes and label classes, respectively. Numbers in the “Split” column are percentages of nodes in training/validation/testing set w.r.t. labeled nodes, respectively
Dataset Nodes $n$ Edges $m$ Features $f$ Classes $N_{c}$ Split
cora (Kipf & Welling, 2017) $2,485$ $12,623$ $1433$ $7$ $0.50/0.25/0.25$
citeseer (Kipf & Welling, 2017) $3,327$ $9,228$ $3703$ $6$ $0.50/0.25/0.25$
pubmed (Kipf & Welling, 2017) $19,717$ $88,648$ $500$ $3$ $0.50/0.25/0.25$
physics (Shchur et al., 2019) $34,493$ $495,924$ $8415$ $5$ $0.50/0.25/0.25$
arxiv (Hu et al., 2020) $169,343$ $2,315,598$ $128$ $40$ $0.54/0.18/0.29$
products (Hu et al., 2020) $2,400,608$ $123,718,024$ $100$ $47$ $0.08/0.02/0.90$
papers100m (Hu et al., 2020) $111,059,956$ $3,228,124,712$ $128$ $172$ $0.78/0.08/0.14$
Appendix C Additional Experiments
C.1 Detailed Experiment Settings
Dataset.
To showcase the scalability of Unifews, we evaluate it on a range of datasets where the backbone models are applicable. Statistics of the datasets can be found in Table 3, where we follow the common settings including the graph preparing and training set splitting pipeline in these benchmarks. In the table, we incorporate self-loop edges and count undirected edges twice to better reflect the propagation overhead.
Evaluation are conducted on a server with 32 Intel Xeon CPUs (2.4GHz), an Nvidia A30 (24GB memory) GPU, and 512GB RAM. Specifically, most iterative sparsification baselines fail to produce results on graphs larger than pubmed due to the out of memory (OOM) error. As they usually involve graph-scale trainable matrix or cache, their memory overhead is typically greater than the raw backbone model, rendering them less applicable to large graphs. In comparison, Unifews performs sparsification on the fly and does not incur additional memory footprint.
Backbone Models.
We select classic GNNs as our backbones, i.e., subject architectures of sparsification methods, due to the consideration that most baselines are only applicable to a couple of classic architectures such as GCN and GAT. In contrast, we demonstrate the generality of Unifews on more backbones including GraphSAGE and GCNII. Iterative backbone models include:
  • GCN (Kipf & Welling, 2017) is the representative message-passing GNN with a diffusion matrix $\bm{T}=\tilde{\bm{A}}$ across all layers.
  • GraphSAGE (Hamilton et al., 2017) performs more specialized message-passing aggregation and is suitable for large graphs.
  • GCNII (Ming et al., 2020) features residual connections and identity mapping. We use $L=32$ for demonstrating the effect of sparsification on deep GNNs.
Baseline Methods.
Our selection of baselines is dependent on their applicable backbones. We mostly utilize their public source codes and retain original implementations. For iterative models, we consider state of the arts in graph and joint compression, which are methods with the capability to produced sparsified graphs with smaller edge sets for both GNN training and inference:
  • Random (RAN) refers to the sparsification method that removes entries with uniform probability based on the specified sparsity.
The graph and weight sparsities are calculated as portions of pruned entries compared to the original matrices. For layer-dependent methods, the average sparsity across all layers is used. For Unifews, the pruning thresholds $\delta_{a}$ and $\delta_{w}$ are determined by Theorem 4.2. Note that as Algorithm 1 naturally preserve diagonal elements in $\hat{\bm{T}}$, a GCN layer with $\eta_{a}=100\%$ Unifews pruning is equivalent to an MLP layer.
C.2 Additional Results on Effectiveness
Extending Figure 3, results of edge and weight sparsification of Unifews and other baselines on more datasets are in Figure 12, while the performance comparison is discussed in Section 5.1. Additionally, we present the experimental results of Unifews on larger datasets and on more backbone architectures in Figure 12 and Figure 12, respectively. No baseline method is available for these settings. It can be observed that Unifews constantly achieves satisfying performance in a large range of sparsity. The effect of graph sparsification on GCNII is relatively unstable, likely because the deeper layers amplify the approximation error.
Variance.
The variance of Unifews is particularly shown in Figure 13 corresponding to Figure 3. Typically, the variance of iterative Unifews models are between the range of 1-3%, which is slightly larger than the backbone model due to the perturbation of edge and weight sparsification.
@@ C.2 Additional Results on Effectiveness
Figure 10: Accuracy of iterative models over citeseer and pubmed, supplementing Figure 3. Columns of “Edge” and “Weight Sparsity” present the average results of models with solely edge and weight sparsification, respectively.
(e) GCN Edge Pruning
(f) GCN Weight Pruning
(g) GCN Edge Pruning
(h) GCN Weight Pruning
Figure 11: Accuracy of iterative Unifews over physics and arxiv, while all baseline methods meet OOM.
(a) GCNII Edge Pruning
(b) GCNII Weight Pruning
(c) GraphSAGE Edge Pruning
(d) GraphSAGE Weight Pruning
Figure 12: Accuracy of iterative Unifews for GCNII and GraphSAGE backbones, while no baseline method is applicable.
(a) GCN Edge Pruning
(b) GCN Weight Pruning
(c) GCN Edge Pruning
(d) GCN Weight Pruning
Figure 13: Accuracy and variance of iterative Unifews over cora.
C.3 Effect of Hyperparameters
Effectiveness of Joint Sparsification.
For effectiveness, the impact of jointly changing $\delta_{a}$ and $\delta_{w}$ is displayed in Figure 14. Intuitively, GCN is more resilient to Unifews sparsification, considering its relatively high redundancy of the wide distribution of entry values. In comparison, GAT is more sensitive to weight removal, that beyond a certain threshold $\delta_{w}$, its accuracy drops significantly. This observation suggests that the value of learnable attention weights in GAT are highly concentrated, and selecting an appropriate sparsity is critical to its performance.